A Doppler ultrasound is used to measure blood flow
Answer #1

a.) The mechanism of ultrasound employs the Doppler effect to give the frequency shift. Initially, the probe acts as the source and the blood cell acts as the object and the sound wave is reflected back from the blood cell and is again received by the probe. In the second case, where the wave is reflected from the blood cell and received by the probe, the blood cell acts as the source and the probe as the receiver. Hence there are two doppler shifts taking place in this mechanism.

b.) Since the blood cell is approaching the probe, there will be a rise in the frequency. So,

the perceived final frequency = initial frequency + frequency shift

f'' = f + 400 = 5 x 106  + 400 = 5000400 Hz

c.) Since the blood cell is moving away from the probe, there will be a drop in the frequncy. So,

perceived frequency = initial frequency - frequency shift

f'' = f - 400 = 5 x 106  - 400 = 4999600 Hz

d.) When the blood cell is moving towards the probe,

the frequency perceived by the blood cell will be f' = f [ ( vwave + vblood ) / vwave ]

when the wave is, in turn, reflected from the blood cell, the blood cell acts as the source now and hence emits a wave of frequency f' (perceived by the blood cell)

So, the final frequency received by the probe is f'' = f' [ vwave / ( vwave - vblood ) ]

f'' = f' [ vwave / ( vwave - vblood ) ] = f [ ( vwave + vblood ) / vwave ][ vwave / ( vwave - vblood ) ]

f'' = f [ ( vwave + vblood ) / ( vwave - vblood ) ]

putting f'' = 5000400 Hz , f = 5000000 Hz , and vwave = 1540 m/s

we get,

5000400 = 5000000 [ (1540 + vblood ) / (1540 - vblood ) ]

7700616000 - 5000400 vblood = 7700000000 + 5000000 vblood

616000 = 10000400 vblood

vblood = 0.0615975361 m/s ~ 0.061 m/s

(Note: if you do the same procedure for the case where blood cell moves away from the probe you'll get a slightly different answer,which is 0.0616024614 m/s. However, this is because the questioner gave the frequency shift for both the cases to be 400 Hz but in reality, they will not be equal. But even then the answer rounded off to three decimals is sufficient for all practical purposes.)